### v2

# RB-SCD: A New Benchmark for Semantic Change Detection of Roads and Bridges in Traffic Scenes

Abstract
With the rapid modernization of urban transportation, the accurate detection of changes such as construction, renovation, and demolition of roads and bridges plays a crucial role in urban planning and traffic management. However, existing methods often struggle to precisely extract semantic change information in such scenarios. A major limitation is the scarcity of fine-grained annotated change detection (CD) datasets in traffic scenes, which hinders the training and optimization of relevant algorithms. To address this, we introduce Road and Bridge Semantic Change Detection (RB-SCD) dataset, a comprehensive benchmark containing 260 pairs of high-resolution remote sensing images. The dataset spans a wide range of geographic regions and includes diverse road and bridge types across more than ten cities in multiple countries, incorporating 11 distinct categories of semantic changes, which collectively enable detailed structural and functional change analysis. Leveraging the rich diversity and complexity of RB-SCD, we further propose a novel framework, Multimodal Frequency-Driven Change Detector (MFDCD), which integrates multiple characteristics of multimodal features in the frequency domain for the first time. Technically, MFDCD consists of two components: Dynamic Frequency Coupler (DFC) and the Textual Frequency Filter (TFF). The DFC deeply couples hierarchical visual features with wavelet-derived frequency components, thereby enhancing the perception of detailed and cross-temporal global structural changes. The TFF converts textual features extracted by the pre-trained CLIP model into the frequency domain via Fourier transform. Then it applies graph-based filtering to extract salient frequency components, which are fused with visual features to enable joint multimodal representation learning. Experimental results show that MFDCD performs well on RB-SCD and three public datasets. The RB-SCD dataset, with its rich semantic annotations, serves as a challenging yet valuable resource for road and bridge change analysis in traffic environments. The code and dataset will be available at https://github.com/DaGuangDaGuang/RB-SCD.
Index Terms:
Change detection (CD), road and bridge semantic change detection dataset, dynamic frequency coupler, textual frequency filter, remote sensing (RS), semantic map guidance.
I Introduction
In the field of remote sensing (RS), change detection (CD) has become an essential task for monitoring environmental, urban, and infrastructure changes [1, 2]. With the acceleration of urbanization and the continuous development of transportation infrastructure, CD in traffic scenes is crucial for traffic planning and facility management. In traffic planning, accurately identifying structural and functional changes in roads and bridges helps urban authorities to assess the current state of transportation networks, forecast future demands, and make data-driven decisions about expansions, detours, and upgrades [3][4][5]. For instance, the emergence of new overpasses, flyovers, or road expansions needs to be timely integrated into transportation maps to ensure up-to-date routing and navigation services. Moreover, detecting illegal road constructions or temporary blockages is also critical for enforcing urban regulations and minimizing congestion. From the perspective of traffic facility management, road and bridge CD enables timely identification of structural deterioration, unauthorized demolitions, or unreported construction activities. This semantic-level understanding is crucial for preventive maintenance, disaster response (e.g., after floods or earthquakes), and resource allocation. Additionally, as smart cities and autonomous transportation systems increasingly rely on accurate and up-to-date digital maps, the need for automated and fine-grained CD becomes even more pressing. Roads and bridges are not only physical assets but also key elements in the digital infrastructure of intelligent transportation systems.
However, as shown in Tab. I, traditional binary change detection (BCD) datasets [6][7][8], etc. primarily focus on building CD and only provide simple information about whether a change has occurred, represented by 0 and 1 for no change and change, respectively. On the other hand, existing semantic change detection (SCD) datasets [9][10][11], etc. mainly focus on changes in land cover, such as building changes and vegetation changes. Changes related to roads and bridges in traffic scenes are either limited in number or completely absent. Most existing road extraction datasets [12][13][14] mainly focus on binary pixel segmentation of roads and cannot be used to focus on road changes. To address this issue, as annotation examples shown in Fig. 1, we propose a new dataset specifically designed for SCD of roads and bridges in traffic scenes: Road and Bridge Semantic Change Detection (RB-SCD) dataset. RB-SCD provides semantic annotations for 11 distinct categories of road and bridge changes, effectively addressing insufficient granularity of information and lack of specificity in detecting road and bridge changes in traffic scenes.
In addition to the RB-SCD dataset, we have also developed a new framework to further facilitate CD performance, the Multimodal Frequency-Driven Change Detector (MFDCD), which integrates multimodal features in the frequency domain. Considering CD methods, traditional approaches often rely on convolutional neural network architectures [15][16][17], which are insufficient to fully exploit the rich spatiotemporal information in images. Some studies [18][19][20] have incorporated transformers into the network for global modeling. However, these methods still face challenges in accurately capturing and effectively distinguishing complex textures and structural changes. Research has also shown that relying solely on single-modal data has inherent limitations. Recently, some methods [21][22][23] have introduced multimodal features to overcome the shortcomings of single modalities, such as combining visual and textual information or integrating lidar and hyperspectral data. However, these approaches often fail to fully exploit the characteristics of multimodal features, focusing only on modeling the commonalities of multimodal data while neglecting the distributional differences among different modalities.
TABLE I: Public remote sensing change detection and road extraction datasets. For image counting, the unit of change detection datasets is pairs, while that of road extraction is pieces/images.
Dataset Resolution (m) Image count Image size (pixels) Type Objects of interest Classes
Change detection
OSCD  [24] 10 24 $600\times 600$ Binary Land cover
S2MLTCP  [25] 10 1520 $600\times 600$ Binary Land cover
ZY3  [26] 5.8 1 $1154\times 740$ Binary Land cover
WV3  [26] 2 2 $1431\times 1431$ Binary Land cover
QB  [26] 2.4 1 $1154\times 740$ Binary Land cover
AICD  [27] 0.5 1000 $800\times 600$ Binary Land cover
LEVIR-CD  [7] 0.5 637 $1024\times 1024$ Binary Buildings
WHU  [6] 0.2 1 $32207\times 15354$ Binary Buildings
CDD  [28] 0.03 - 1 16000 $256\times 256$ Binary Buildings
SYSU  [8] 0.5 20000 $256\times 256$ Binary Buildings
BDD  [29] 1.0 1147 $250\times 250$ Semantic Buildings 1
Landsat-SCD  [10] 30 8468 $416\times 416$ Semantic Land cover 4
HRSCD  [9] 0.5 291 $10000\times 10000$ Semantic Land cover 5
SECOND  [11] 0.5 - 3 4662 $512\times 512$ Semantic Land cover 6
Hi-UCD  [30] 0.1 1293 $1024\times 1024$ Semantic Land cover 9
Road extraction
CHN6-CUG  [12] 0.5 4511 $512\times 512$ Binary Road
Massachusetts  [13] 1.5 1000 $1500\times 1500$ Binary Road
Google Earth  [14] 1.2 224 $512\times 512$ Binary Road
RB-SCD (Ours) 0.59 260 $270\times 534$ $7215\times 4366$ Semantic Roads and bridges 8
Figure 1: Annotation examples on RB-SCD dataset. $T_{1}$ and $T_{2}$ represent the pre- and post-change images, respectively. Label is the ground truth.
Figure 2: Visualization of visual and textual frequency spectrum.
In this work, as shown in Fig. 2, we observe that the decomposition and utilization of features in the frequency domain remain underexplored, particularly the fine-grained change information embedded in the high-frequency visual features, the large-scale distribution information in the low-frequency features, and the frequency-domain representation of textual features. The potential patterns of multimodal information in the frequency domain have not been adequately studied. Therefore, we make the first attempt to develop a new architecture that maps image and textual features into different types of frequency domains based on their characteristics, enabling multimodal frequency-domain modeling for CD.
Our MFDCD consists of two main components: Dynamic Frequency Coupler (DFC) and the Textual Frequency Filter (TFF). DFC decomposes bi-temporal features into high- and low-frequency components using discrete wavelet transform (DWT), and introduces an Adaptive Sparse Frequency Fusion (ASFF) block for the combination of high-frequency information and low-level features. During the early stages of training, ASFF enhances the sparsity of high-frequency features, focusing on significant change areas. In the later stages, the sparsity constraint is reduced to enhance the expressive power of the model, ultimately increasing the attention to key change areas while suppressing high-frequency noise interference. In this way, edge and texture information in low-level features can not only be gradually captured through convolution operations but also dynamically weighted based on the localized changes in refined high-frequency information. For low-frequency features, Bidirectional Temporal Frequency Fusion (BTFF) block is introduced. Since CD involves not only capturing local changes but also understanding the relative changes between different time points, simply combining the high-level features at a single time point with low-frequency features could cause the model to overly rely on the local structural information at that moment, neglecting the cross-temporal changes. The BTFF block addresses this by combining high-level features with low-frequency features across time points, enabling the model to understand the trend of changes and global structural changes over time.
TFF introduces a pre-trained CLIP [31] model to obtain textual information, which is further expanded to get text descriptors. Text features contain semantic information but lack spatial features, and discrete Fourier transform (DFT) can convert them into frequency domain information, making it easier to align with the frequency information of image features. By constructing graph filters to aggregate key information, the processed textual features are deeply integrated with visual features to provide multimodal support for SCD.
Our contributions are as follows:
  • A comprehensive benchmark for SCD in traffic scenarios. We construct a new dataset named RB-SCD, which focuses on SCD of roads and bridges. It contains 260 large-scale high-resolution RS image pairs, covering over ten cities across multiple countries and annotating 11 typical semantic change types. With its wide geographic diversity and complex traffic infrastructures, RB-SCD supports fine-grained structural and functional change analysis for transportation environments.
  • A novel multimodal frequency-driven CD framework. We propose Multimodal Frequency-Driven Change Detector (MFDCD), the first framework to introduce frequency-domain multimodal feature integration for CD. It consists of two main modules: (1) Dynamic Frequency Coupler (DFC), which fuses hierarchical visual features with high- and low-frequency components obtained via wavelet transform to enhance cross-temporal structural change perception; (2) Textual Frequency Filter (TFF), which transforms textual features extracted by a pre-trained CLIP model into the frequency domain using Fourier transform, applies graph-based filtering to extract salient components, and integrates them with visual features for collaborative representation learning.
  • Comprehensive experiments on multiple benchmarks. Extensive experiments on the proposed RB-SCD and three public CD datasets demonstrate that MFDCD achieves superior performance and generalization ability in identifying complex semantic changes in traffic scenes.
II Related Work
II-A CD Datasets and Methods
Existing datasets, as shown in Tab. I, such as  [6][7][8][28], are primarily designed for BCD tasks and cannot provide specific semantic change information. For SCD, datasets like  [9][10][11] have established some benchmarks. However, these datasets mainly focus on changes in land cover, with little attention paid to the changes in roads and bridges in traffic scenes. In contrast, our RB-SCD dataset is a high-resolution, fine-grained SCD dataset specifically curated and annotated for road and bridge changes in traffic scenes. It captures a wide range of semantic change types and includes diverse geographical contexts, enabling precise and meaningful evaluation of CD algorithms in real-world urban transportation scenarios. By addressing both the lack of semantic detail and the absence of domain specificity in existing datasets, RB-SCD provides a valuable and targeted benchmark for infrastructure monitoring, intelligent transportation analysis, and traffic scene understanding.
Figure 3: (a) Statistical distribution of pixels for semantic change categories. (b) Train set. (c) Test set.
TABLE II: Specific subcategories in RB-SCD dataset.
Categories Description
Road street, highway, dirt road, walkway, railway, etc
Bridge
beam bridge, arch bridge, cable-stayed bridge,
suspension bridge, etc.
Bare land
wasteland, sand patches area, unpaved open
grounds, road or railway construction zones, etc.
Grass
urban lawn, green belt, wetland grass, riparian
grass zone, meadow etc.
Water
ocean, lake, river, reservoir, pond, etc.
Farmland agricultural land (with and without crops)
Vegetation forest, shrub, urban vegetation buffer, etc.
Building
residential building, commercial building,
industrial building, under-construction building
(with clear structure), etc.
In recent years, deep learning has significantly improved the accuracy of CD through its ability to extract relevant features from raw data. Changemask [32] decouples SCD into temporal semantic segmentation and BCD, and combines these two tasks into a unified encoder-transformer-decoder framework. MDENet [33] uses MDCM to capture object features in frequency and spatial domains across times, and MFGatt to reduce style impact, focusing on object location. JFRNet [34] introduces a three-branch network that takes joint bi-temporal features as unified input, enabling each branch to perceive the semantic relevance of the two time points without any additional interaction operations. DecoderFusion [35] enhances CD decoding features by leveraging the feature differences between two semantic segmentation decoding branches. CdSC [36] introduces a cross-differential semantic consistency network that explores the differences in bi-temporal features while maintaining their semantic consistency. HGINet [37] builds a multi-level perception aggregation network with a pyramid architecture to extract features that distinguish different categories across multiple levels. However, these methods only consider single-modal data, i.e., images, while neglecting the rich semantic information embedded in multimodal data, thus encountering a performance bottleneck.
II-B Multimodal Learning
Recent studies have shown that multimodal information is beneficial for visual image analysis. LAVT [21] and VLT [38] integrate textual features into visual features through language encoders. Zhang et al. [22] propose a convolutional neural network method with an attention mechanism to enhance feature extraction for lidar data. Their approach leverages the elevation information from lidar data and the spatial-spectral information of hyperspectral data for multimodal feature fusion. MedCLIP [39] uses text prompts to generate segmentation for clinical scans, adopting a novel decoupled hard-negative noise contrastive estimation loss to fine-tune the BiomedCLIP model. ChangeCLIP [23] designs a novel CD framework that utilizes semantic information from image-text pairs and introduce an innovative differential feature compensation module to capture fine-grained semantic changes between them.
II-C Frequency Transform in Deep Learning
In image processing, the combination of CNN and frequency domain information has been widely applied in tasks such as image restoration, super-resolution, and semantic segmentation. W-Unet [40] replaces down-sampling operations with DWT to reduce the impact of image noise while avoiding information loss. FFT-ReLU  [41] performs image deblurring tasks in the frequency domain based on fast Fourier transform. ARFFT [42] proposes a spatial-frequency fusion block to enhance the representation capability of Transformers and extend the receptive field across the entire image. XNet [43] utilizes wavelet transforms to transfer visual information to the frequency domain, leveraging high- and low-frequency information to design a framework supporting both fully supervised and semi-supervised tasks. Given the characteristics of RS images, such as high resolution and susceptibility to spectral variations caused by seasonal changes, using only frequency domain features often results in the loss of critical spatial information. Therefore, in this paper, we propose a novel architecture that models multimodal features in the frequency domain, effectively balancing the advantages of visual features, textual features, and frequency domain features, achieving promising results.
Figure 4: Examples of dataset distribution spanning multiple representative geographical areas.
Figure 5: Examples of road and bridge types, where the first two columns are roads, and the last three columns are bridges.
III RB-SCD Dataset
III-A Data Annotation and Analysis
Rich Semantic Annotations: As shown in Fig. 1 and Fig. 3, unlike traditional BCD datasets, we have defined eight major categories (with specific subcategories detailed in Tab. II), including road, bridge, bare land, building, farmland, water, grass and vegetation. The RB-SCD dataset provides detailed semantic change annotations for 11 distinct types of semantic changes and a background category across these eight categories.
Figure 6: Overview of our proposed Multimodal Frequency-Driven Change Detector (MFDCD). Firstly, the Dynamic Frequency Coupler (DFC) couples hierarchical visual features extracted by ResNet50 with high- and low-frequency components obtained via wavelet transform to enhance cross-temporal structural change perception. Meanwhile, the Textual Frequency Filter (TFF) transforms textual features extracted by a pre-trained CLIP model into the frequency domain using Fourier transform, applies graph-based filtering to extract salient components, and integrates them with visual features for collaborative representation learning. Finally, encoder features are fed into the decoder to obtain the final semantic change detection results.
Changes such as road widening, new road construction, and bridge construction or modification are reflected through the corresponding semantic change types, supporting in-depth exploration of the intrinsic mechanisms of traffic scenario evolution. It is worth noting that some changes, such as bridges directly turning into roads, are rarely observed in reality. Additionally, other types of changes, such as those between bridges and some other categories, occur in only a few samples, and we set their pixel number to 0. All annotations are carried out by a professional team in RS and transportation, and underwent multiple rounds of rigorous review and quality control to ensure annotation accuracy and semantic consistency.
Diverse Geographic Coverage: As shown in Fig. 4, following the principles in [44], we collect 260 pairs of RS images with resolutions ranging from 270 × 534 to 7,215 × 4,366 (0.59 m/pixel), which can be cropped into 5,016 pairs of 256 × 256 patches, including 4,015 pairs for the training set and 1,001 pairs for the testing set. The dataset integrates satellite images obtained primarily from Google Earth, covering various regions worldwide, including Chinese cities such as Hefei, Xi’an, Chongqing and Guangzhou, as well as representative areas from Japan, Singapore, Nigeria, Italy, America, and Indonesia, etc. This extensive geographic coverage ensures that the dataset encompasses traffic scenes under varying climatic conditions, geological structures, and urban planning contexts, thereby guaranteeing its diversity and representativeness.
Multiple Road and Bridge Types: As shown in Fig. 5, the dataset encompasses various types of roads and bridges from different geographical regions and urban environments. The road categories include concrete and asphalt roads used for urban main roads, secondary roads, residential streets, and highways, comprehensively reflecting the common transportation routes found in cities. Additionally, the dataset contains elevated roads designed to alleviate congestion and improve traffic efficiency, which are typically found in densely populated or high-traffic urban areas. In terms of bridge structures, the dataset includes large cable-stayed bridges that span wide rivers or deep valleys, as well as girder bridges and pedestrian bridges used for short- to medium-span connections and pedestrian access. The dataset also includes other types of specialized roads and bridges, demonstrating a variety of engineering styles and applications. This richness in road and bridge types enables the dataset to represent complex traffic environments and provides robust support for training CD models across multiple scales and diverse scenarios.
III-B Label Imbalance Analysis and Challenges
As shown in Fig. 3, our RB-SCD dataset exhibits a certain degree of sample imbalance In terms of semantic changes, there is an imbalance in the types of changes in the dataset, especially for fine-grained CD tasks. For some types of changes (e.g., road widening, bridge construction), there are more samples, while for other types of changes (e.g., road repair, demolition), there are fewer samples. This imbalance may lead to the model overfitting to the more common change types, while neglecting the rarer ones. Additionally, the imbalance between change and no-change samples is another issue, particularly for CD tasks. The no-change samples are relatively more abundant, which may lead to a decrease in the model’s sensitivity when processing change samples. To improve the accuracy and robustness of the model, we will further collect more samples of different types of changes in future work and employ techniques such as data augmentation and sample weighting to alleviate the model’s difficulty in detecting rare changes.
IV Methodology
IV-A Overview
As shown in Fig. 6, our MFDCD mainly includes two components: (1) DFC module effectively couples high-level and low-level visual features with high-frequency and low-frequency spectral features. (2) TFF module enhances the text information obtained through zero-shot inference, explores the representation of text features in the frequency domain, and promotes the interaction and fusion of multimodal data. The details are as follows.
IV-B DFC Module
Given the bi-temporal images $T_{1}$ and $T_{2}$, ResNet50 [45] is used as the backbone network to extract the corresponding multi-scale features $F_{i}^{t}(i\in\left\{1,2,3,4\right\},t\in\left\{1,2)\right\}$. The features from the first two layers of the network are considered as low-level features, while the features from the last two layers are considered as high-level features. The low-frequency and high-frequency components of these features are then obtained through wavelet transform, which allows for the separation of coarse and fine-grained information in both spatial and frequency domains.
Adaptive Sparse Frequency Fusion: Low-level features are rich in detailed information, while high-frequency features contain information about edges, details, and abrupt changes, which naturally correspond to low-level features. However, high-frequency features are sensitive to noise and may amplify the influence of irrelevant regions. As shown in the upper-right corner of Fig. 6, the ASFF block applies a novel sparsity constraint to the high-frequency features, which is dynamically adjusted during training iterations. By combining the sparsity-constrained high-frequency features with low-level features, both the image’s detailed information (such as edges and textures) and local structural information are preserved. This combination makes the network more responsive to CD, especially in areas with detailed changes. The process is as follows:
$LL_{i}^{t},HL_{i}^{t},LH_{i}^{t},HH_{i}^{t}=\mathcal{W}(F_{i}^{t}),$ (1)
where $t\in\left\{1,2\right\}$, $i\in\left\{1,2,3,4\right\}$, $\mathcal{W}(\cdot)$ represents the discrete wavelet transform. $LL_{i}^{t}$ is the low-frequency band, $HL_{i}^{t}$, $LH_{i}^{t}$, and $HH_{i}^{t}$ signifying the high-frequency bands.
$H_{i}^{t}=Conv(Cat(HL_{i}^{t},LH_{i}^{t},HH_{i}^{t})),$ (2)
$\widetilde{H_{i}^{t}}=\mathcal{S}(H_{i}^{t},\lambda(p),g)=\begin{Bmatrix}H_{i}^{t},&if\left|H_{i}^{t}\right|>g\cdot\lambda(p)\\ 0,&otherwise\end{Bmatrix},$ (3)
$\lambda(p)=\lambda_{0}\cdot e^{-\gamma p},$ (4)
$P_{i}^{t}=Conv(Cat(Up(\widetilde{H_{i}^{t}}),F_{i}^{t})),$ (5)
where $t\in\left\{1,2\right\}$, $i\in\left\{1,2\right\}$. $\mathcal{S(\cdot)}$ denotes the sparsity constraint function, which is used to apply the sparsity operation, $\lambda(p)$ is the dynamic sparsity strength, $\lambda_{0}$ is the initial sparsity strength, $\gamma$ is the sparsity decay rate, $p$ is the current training iteration, and $g$ is a constant controlling the sparsity strength. $H_{i}^{t}$ is the original fused high-frequency feature and $\widetilde{H_{i}^{t}}$ is the high-frequency feature after applying the sparsity constraint. $P_{i}^{t}$ is the enhanced feature that combines high-frequency and low-level information. $Conv(\cdot)$ and $Up(\cdot)$ are convolution and bilinear upsampling operation, respectively.
Bidirectional Temporal Frequency Fusion: As shown in the lower-right corner of Fig. 6. Take the low-frequency features of time $T_{2}$ and the high-level features of time $T_{1}$ as an example, high-level features at time $T_{1}$ typically contain the semantic hierarchy of the image, enabling the expression of global structural changes in the image, such as road widening and bridge construction. Low-frequency features at time $T_{2}$ mainly focus on smooth regions or large-scale change information in the image. By combining high-level features from $T_{1}$ with low-frequency features from $T_{2}$, the network can capture global structural changes across time points while maintaining sensitivity to large-scale, smooth region changes. This models the bidirectional relationship of changes: local changes at $T_{1}$ may lead to semantic-level changes at $T_{2}$, while detailed changes at $T_{2}$ may arise from regional distribution adjustments at $T_{1}$. The details of BTFF block are as follows:
$P_{i}^{1}=Conv(Cat[Up(LL_{i}^{2}),F_{i}^{1}]),$ (6)
$P_{i}^{2}=Conv(Cat[Up(LL_{i}^{1}),F_{i}^{2}]),$ (7)
where $i\in\left\{3,4\right\}$, $P_{i}^{t}$ is the enhanced feature that combines low-frequency information and high-level information.
IV-C TFF Module
As shown in Fig. 6, we leverage the powerful zero-shot inference capability of the CLIP model to obtain textual information from RS images for multimodal modeling. For the categories inferred by the CLIP model, we select the top five categories with the highest probability values. For the foreground features, the text description $T_{d}$ is designed in the format “Remote sensing image at time $\left\{period\right\}$ has a $\left\{value\right\}$ probability of being the $\left\{key\right\}$ ,” where $\left\{period\right\}$ represents either $T_{1}$ or $T_{2}$, $\left\{value\right\}$ is the probability value of the corresponding category and $\left\{key\right\}$ represents the category itself. $E_{T}$ is a pre-trained CLIP text encoder that encodes the text embedding. Considering that discrete Fourier transform can help to reveal high-order semantic patterns in text, especially by identifying more complex semantic information through the descriptors used at different temporal points ($T_{1}$, $T_{2}$). Through Fourier transform, periodic or repetitive structures in the text can be extracted, which helps identify commonalities and differences in changes. We first use Fourier transform to convert the text embeddings from the text domain to the frequency domain:
$\mathbf{T}_{freq}=\mathcal{F}(E_{T}(T_{d})),$ (8)
where $\mathcal{F}(\cdot)$ is the discrete Fourier transform, $\mathbf{T}_{freq}\in\mathbb{C}^{B\times N\times C}$, $B$ represents the batch size, $N$ represents the number of frequency components, $C$ represents the dimension of each frequency component. We introduce a graph filter bank $K=[k_{1},k_{2},...k_{k}]$ for the textual features, which propagate information between the nodes of the graph to filter and refine the frequency information, thereby obtaining salient features related to the changes. For the frequency feature $\mathbf{T}_{freq}$, we split it into real and imaginary parts and concatenate them to construct node features. The feature matrix $\mathbf{X}\in\mathbb{R}^{B\times N\times 2C}$ is the collection of features for all nodes, which allows us to convert frequency domain information into node features for the graph. Assuming that each frequency component is connected to every other frequency component (fully connected graph), the edge set $\mathbf{E}\in\mathbb{Z}^{2\times\frac{N(N-1)}{2}}$ of the graph can be constructed by combining each pair of nodes:
$\mathbf{X}=Cat[\mathfrak{R}(\mathbf{T}_{freq}),\mathfrak{I}(\mathbf{T}_{freq})],$ (9)
$\mathbf{E}=\left\{(i,j)\mid 1\leq i\neq j\leq N\right\},$ (10)
$\mathbf{E}^{(k+1)}=\sigma(\hat{\mathbf{A}}\mathbf{E}^{(k)}\mathbf{W}_{k}),$ (11)
$\mathbf{E}^{f}=\sum_{k=1}^{k}\mathbf{E}^{(k)},$ (12)
where $\mathfrak{R}(\mathbf{T}_{freq})$ and $\mathfrak{I}(\mathbf{T}_{freq})$ represent the real and imaginary parts of the frequency domain features, respectively. $\mathbf{E}^{(k)}$ represents the node features after the $k$-th filter, $\hat{\mathbf{A}}$ is the normalized adjacency matrix, which represents the relationships between nodes, $\mathbf{W}_{k}$ is the weight matrix of the graph convolution filter at the $k$-th layer, and $\sigma$ is the activation function, $\mathbf{E}^{f}$ is the weighted sum of the outputs from multiple filters. We use the text-image encoder $E_{VT}$ of the CLIP model to fuse $\mathbf{E}^{f}$ and the image features. After subtracting the results from the bi-temporal phases, the output is fed into the decoder [46] to obtain our CD results.
IV-D Frequency Transform Function Analysis
Why use discrete wavelet transform for image features and discrete Fourier transform for textual features? The advantage of wavelet transform lies in its ability to analyze images at different scales and locations, making it particularly suitable for multi-level, multi-scale image analysis tasks. As shown in Fig. 2, in remote sensing image CD, wavelet transform helps to extract high-frequency features (e.g., detailed changes) and low-frequency features (e.g., overall structure), aiding in the capture of changes at different spatiotemporal scales in the image. Considering that text is typically discrete sequential data, its changes are often global or frequency-related. Therefore, on the other hand, the Fourier transform can effectively reveal periodic semantic patterns in the text, which contributes to identify patterns and trends in text changes.
V Experiments and Results
V-A Public CD Datasets
In addition to using our proposed dataset, as shown in Tab. III, we have employed three public CD datasets to evaluate the robustness of our model. The LEVIR-CD [7] dataset consists of 637 pairs of RS images with a resolution of 1,024 × 1,024. It mainly focuses on building area changes, such as building growth. The WHU-CD [6] dataset comprises a pair of high-resolution RS images with a size of 32,507 × 15,354 pixels. It covers areas before and after an earthquake and primarily includes building renovations and additions. The SYSU-CD [8] dataset is a collection of 20,000 pairs of high-resolution aerial images, each with a size of 256 × 256 pixels. The images are captured in the Hong Kong area and include urban construction, suburban expansion, and vegetation changes, etc.
TABLE III: Public CD datasets used in the experiments.
Datasets Train/ Val/ Test Image type Space resolution
LEVIR-CD 7120/ 1024/ 2048 Satellite images 0.5 m/pixel
SYSU-CD 12000/ 4000/ 4000 Aerial images 0.5 m/pixel
WHU-CD 5947/ 743/ 744 Aerial images 0.2 m/pixel
TABLE IV: Comparison of quantitative performance for semantic change detection on RB-SCD dataset. All metrics are expressed as percentages (%). Results highlighted in red indicate the best performance, while those in blue denote the second-best performance.
Class HGINet [37] CdSC [36] DEFO [35] ChangeCLIP [23] Ours
IoU F1 Rec Pre IoU F1 Rec Pre IoU F1 Rec Pre IoU F1 Rec Pre IoU F1 Rec Pre
background 97.76 98.87 99.14 98.60 98.22 99.1099.45 98.75 98.18 99.08 98.27 98.90 98.11 99.05 99.37 98.73 98.28 99.13 99.34 98.93
bare land ->road 45.24 62.30 61.10 63.54 48.42 65.25 63.40 67.21 50.93 67.49 66.0469.00 51.75 68.21 64.87 64.87 52.76 69.0869.4968.67
water ->bridge 80.20 89.01 85.60 92.71 85.33 92.09 93.32 90.88 83.10 90.77 91.38 90.17 85.77 92.34 88.22 96.87 86.83 92.9594.28 91.67
water->road 58.09 73.49 76.88 70.39 68.64 81.49 74.82 89.26 62.78 77.14 76.21 78.88 68.11 81.03 76.23 86.47 71.68 83.51 79.6387.78
road->bare land 33.83 50.55 40.13 68.29 38.33 55.42 44.58 73.22 38.14 55.22 48.02 64.94 40.57 57.7247.0674.63 39.48 56.61 43.88 79.73
road->vegetation 10.69 19.32 12.56 41.82 13.65 24.03 15.25 56.57 17.02 29.0922.47 41.25 16.37 28.13 19.35 51.49 20.31 33.7621.8374.39
vegetation->road 52.61 68.94 63.10 75.98 57.93 73.36 69.89 77.19 57.84 73.29 72.30 74.41 57.10 72.69 68.55 77.37 59.35 74.49 72.38 76.72
building->road 36.56 55.57 43.44 69.78 36.10 53.05 41.28 74.18 39.83 56.97 51.63 63.54 40.46 57.61 44.69 81.04 45.44 62.49 53.0675.98
farmland->road 63.36 73.55 77.80 77.34 63.80 77.90 78.78 86.61 65.38 79.07 75.87 82.55 65.02 78.80 81.30 76.45 68.40 81.24 79.21 83.38
grass->road 36.55 53.54 39.09 84.94 36.66 53.65 46.45 64.48 40.60 57.7548.26 71.89 36.55 53.54 39.09 84.94 42.49 59.6447.24 80.86
road->grass 22.44 36.65 32.38 42.21 15.53 26.89 17.41 58.97 18.41 31.10 22.75 49.14 21.76 35.7524.2168.26 21.65 35.59 22.86 80.30
bridge->water 46.51 63.49 49.92 87.19 79.36 82.60 74.25 93.06 83.40 90.9593.45 88.57 77.86 87.55 90.44 84.85 83.58 91.05 90.34 91.78
mIoU 48.65 52.75 54.63 54.95 57.52
TABLE V: Performance of the RS foundation model LWGANet on RB-SCD dataset. All metrics are expressed as percentages (%).
Class LWGANet-L1 [47] LWGANet-L2 [47]
IoU F1 Pre Rec IoU F1 Pre Rec
background 98.30 99.14 98.84 99.44 98.27 99.13 98.8 99.45
bare land ->road 54.17 70.27 73.28 67.50 53.17 69.42 74.46 65.02
water ->bridge 87.29 93.21 92.52 93.92 85.85 92.39 90.95 93.87
water->road 69.53 82.03 82.51 81.55 63.42 77.62 83.84 72.25
road-> bare land 38.54 55.64 78.94 42.95 38.06 55.14 74.67 43.71
road->vegetation 15.54 26.89 55.44 17.75 13.70 24.10 68.9 14.60
vegetation->road 59.14 74.32 75.55 73.13 60.10 75.08 78.81 71.68
building->road 42.10 59.25 81.97 46.40 42.27 59.42 76.50 48.57
farmland->road 69.94 82.31 86.16 78.79 71.04 83.07 83.96 82.2
grass->road 28.98 44.93 81.75 30.98 40.78 57.93 81.93 44.81
road->grass 22.78 37.11 83.44 23.86 21.87 35.89 84.47 22.79
bridge->water 75.10 85.78 89.44 82.41 80.01 88.90 89.56 88.24
mIoU 55.12 55.71
TABLE VI: Comparison of quantitative performance for binary change detection on RB-SCD dataset. All metrics are expressed as percentages (%). Results highlighted in red indicate the best performance, while those in blue denote the second-best performance.
Model IoU F1 Rec Pre
FC-diff [15] 57.81 73.26 69.63 77.30
FC-ef [15] 59.05 74.26 69.05 80.31
FC-cat [15] 60.34 75.26 71.99 78.85
SNUNet [16] 62.90 77.22 72.47 82.64
BIT [18] 62.21 76.71 70.97 83.69
ChangeFormer [19] 61.88 76.45 69.69 84.66
ICIF [17] 61.80 76.39 71.93 81.43
TFI-GR [48] 64.72 78.58 73.78 84.05
DMINet [20] 62.99 77.29 72.14 83.24
USSFC-Net [49] 62.55 76.96 75.07 78.95
ELGC-Net [50] 65.53 79.18 73.99 85.15
Ours 65.83 79.39 74.13 85.47
TABLE VII: Comparison of quantitative performance of various CD techniques on public datasets. All metrics are expressed as percentages (%). Results highlighted in red indicate the best performance, while those in blue denote the second-best performance. Results marked with * are obtained from the original papers.
Model LEVIR-CD [7] SYSU-CD [8] WHU-CD [6]
IoU  F1  Rec  Pre IoU  F1  Rec  Pre IoU  F1  Rec  Pre
FC-diff [15] 75.18 85.83 84.59 87.12 42.05 59.20 46.72 80.80 69.48 81.99 78.90 89.14
FC-ef [15] 71.80 83.59 82.27 84.94 55.83 71.65 65.41 79.22 58.53 71.65 65.41 79.22
FC-cat [15] 76.98 86.99 85.64 88.38 60.68 75.53 71.38 80.21 72.24 83.87 84.22 83.53
SNUNet [16] 82.20 90.23 88.97 91.54 66.73 80.04 79.79 80.30 76.02 86.38 88.95 83.95
BIT [18] 82.06 90.15 88.98 91.34 62.50 76.93 72.50 81.93 80.40 89.13 89.69 88.58
ChangeFormer [19] 81.93 90.07 87.38 92.92 60.46 75.36 71.06 80.22 83.88 91.23 90.76 91.71
ICIF [17] 79.86 88.80 86.85 90.84 61.38 76.07 73.31 91.94 78.27 87.81 84.06 91.96
TFI-GR [48] 83.03 90.73 89.53 91.95 71.38 83.30 81.16 85.56 88.55 93.92 92.65 95.24
DMINet [20] 83.14 90.79 88.89 92.78 68.20 81.10 76.82 85.88 79.53 88.60 91.24 86.10
USSFC-Net [49] 82.54 90.44 91.82 89.09 65.75 79.34 81.52 77.27 83.92 91.26 93.20 89.40
M-Swin * [51] 83.57 91.05 90.05 92.07 -   -   -   - 86.09 92.94 91.37 94.56
ELGC-Net [50] 83.52 91.02 89.78 92.30 65.34 79.03 77.26 80.89 75.56 86.08 82.15 90.40
MDENet * [33] 83.63 91.0992.68 89.55 -   -   -   - 87.74 93.47 95.49 91.54
Ours 84.69 91.7192.31 91.11 72.63 84.15 85.60 82.74 90.43 94.97 96.44 93.55
V-B Implementation Details
Our experiments are conducted on a single NVIDIA GeForce RTX 3090 GPU. The AdamW [52] is employed, with a learning rate of 0.0001 and a weight decay parameter of 0.01. Cross entropy loss is used as our loss function, and the maximum number of iterations is set to 40,000. For data augmentation, RandomFlip and PhotoMetricDistortion are applied. The constant $g$ in the ASFF block is set to 0.1, the decay rate $\gamma$ is set to 0.0001, and $\lambda_{0}$ is set to 1. To evaluate the performance of our experiments, we adopt four widely used metrics: recall (Rec), precision (Pre), F1-score (F1) and intersection over union (IoU), where IoU and F1 are the primary metrics, while the others are auxiliary metrics. The following are formulas for defining the evaluation metrics as mentioned above.
$Rec=TP/(TP+FN),$ (13)
$Pre=TP/(TP+FP),$ (14)
$F1=2/(Rec^{-1}+Pre^{-1}),$ (15)
$IoU=TP/(TP+FP+FN),$ (16)
where TP, TN, FP, and FN denote the counts of true positives, true negatives, false positives, and false negatives, respectively.
V-C Quantitative Assessment Results
Results for Semantic Change Detection on RB-SCD Dataset: Tab. IV presents the comparative results of our method and several advanced SCD methods across multiple evaluation metrics. As shown in the table, we have defined 11 types of semantic changes along with one background category, aiming to comprehensively cover the diverse real-world scenarios of road and bridge changes. These change types include both relatively straightforward categories—such as water to bridge and farmland to road, which typically involve simple backgrounds or abundant training samples—and more challenging categories, such as road to vegetation and grass to road, which are characterized by complex backgrounds, ambiguous boundaries, and limited sample sizes. Across this diverse range of semantic changes, our method achieves significant advantages in both the IoU and F1 metrics. Specifically, for the easily detectable categories—water to bridge, bridge to water, water to road, and farmland to road—our model attains IoU scores of 86.83%, 83.58%, 71.68%, and 68.40%, respectively, significantly outperforming other comparison methods. These results highlight the strong capability of our model in semantic discrimination and spatial reasoning. More importantly, our method also performs excellently on several challenging categories, including road to vegetation, road to grass, road to bare land, and grass to road, which suffer from sample scarcity and semantic ambiguity. These types of changes are often affected by similar surface textures, seasonal variability, and complex contextual information. Traditional methods are prone to false positives and missed detections under such conditions, whereas our model consistently demonstrates strong robustness and generalization ability by accurately identifying these subtle semantic transformations.
Furthermore, we evaluate the performance of the latest RS foundation model, LWGANet [47], on our proposed dataset (as shown in Tab. V). The results further underscore the challenging nature of our dataset. By incorporating fine-grained and practically meaningful semantic change categories in traffic scenarios, our dataset provides a more rigorous and instructive benchmark for training and evaluating CD models in real-world applications.
Results for Binary Change Detection on RB-SCD Dataset: As shown in Tab. VI, in the BCD task based on the RB-SCD dataset, different models exhibit varying performance across different evaluation metrics. In terms of the IoU metric, our model ranks first with a score of 65.83%, slightly outperforming ELGC-Net [50], which achieves 65.53%. This indicates that MFDCD demonstrates the best performance in terms of overlap between the predicted and ground truth change regions, enabling more precise localization of change areas. Regarding the F1, our MFDCD model also leads with a score of 79.39%, followed closely by ELGC-Net at 79.18%. Since the F1 is a harmonic mean of Pre and Rec, this result suggests that MFDCD achieves an excellent balance between detection accuracy and completeness. For recall, USSFC-Net [49] achieves the highest score of 75.07% among all models, while our model follows closely with a recall of 74.13%, ranking second. This demonstrates that USSFC-Net has a strong ability to detect all change areas, and our model also exhibits competitive performance in capturing changed regions. In terms of precision, our model outperforms others with a score of 85.47%, while ELGC-Net follows at 85.15%. This implies that our model has a lower false positive rate when identifying change areas. Overall, the results of the BCD task further indicate that existing methods still have room for improvement when applied to our RB-SCD dataset.
Results on Public CD Datasets: To further prove the validity of our model, we also compare our proposed MFDCD with thirteen state-of-the-art CD techniques. To ensure a fair comparison, we reproduce these methods using publicly available code with default parameters. Tab. VII presents the quantitative evaluations of various methods on three public CD datasets, including IoU, F1, Rec, and Pre metrics. As observed in Tab. VII, our method achieves outstanding results across all three publicly available datasets. Specifically, on the WHU-CD dataset, our approach outperforms the second-best method by 2.69% and 1.5% in IoU and F1, respectively. On the SYSU-CD dataset, it achieves improvements of 4.43% and 3.05%, while on the LEVIR-CD dataset, it surpasses the second-best method by 1.06% and 0.62%. These results further demonstrate the effectiveness and robustness of our method.
Figure 7: Qualitative visualization for semantic change detection on RB-SCD dataset. (a) $T_{1}$ images. (b) $T_{2}$ images. (c) HGINet. (d) CdSC. (e) DEFO. (f) ChangeCLIP. (g) Ours. (h) Label.
Figure 8: Qualitative visualization for binary change detection on RB-SCD dataset. (a) $T_{1}$ images. (b) $T_{2}$ images. (c) BIT. (d) ChangeFormer. (e) ICIF. (f) TFI-GR. (g) DMINet. (h) USSFC-Net. (i) ELGC-Net. (j) Ours. (k) Label. The rendered colors represent true positives (white), false positives (red), true negatives (black), and false negatives (blue).
Figure 9: Heatmaps of our proposed modules. (a) and (g) Input images. (b) and (h) Low-level features from the backbone network. (c) and (i) High-level features from the backbone network. (d) and (j) Feature maps after undergoing the ASFF block. (e) and (k) Feature maps after undergoing the BTFF block. (f) and (l) Feature maps after undergoing the TFF module.
V-D Quantitative Analysis.
As shown in Fig. 7, we compared the visualized results of the proposed MFDCD with other state-of-the-art methods on RB-SCD dataset. From the figure, it can be observed that our method outperforms others in identifying small semantic change regions. For instance, as shown in the third row, our method better distinguishes between changes from grass to road and from vegetation to road—two types of changes with similar characteristics. Regarding pseudo-changes caused by background composition, such as in the fifth row, where a boat passing under a bridge with its bow color similar to the bridge’s edge color, both DEFO [35] and ChangeCLIP [23] made incorrect predictions, while our method accurately detected the changes. Additionally, the proposed dataset presents significant challenges. For example, in the first row, all methods struggle to detect semantic changes under road shadows. As shown in the second row, farmland in autumn and winter appears visually similar to bare land, and in the fourth row, farmland in spring and summer shares similar colors with water, resulting in low inter-class differences. In such challenging scenes, other methods show many false detections, while our method performs slightly better but still leaves room for improvement.
When comparing the BCD results of each method with the ground truth in Fig. 8, it can be observed that for the unchanged areas (black regions representing true negatives), most models show relatively accurate results, with fewer false positives. In contrast, for the changed areas, the performance varies among models. Some models exhibit noticeable missed detections (blue regions representing false negatives), indicating they fail to identify all actual changes. Many models are capable of detecting building—related changes, which are common in the dataset. For example, in areas where buildings are newly constructed or modified, most models can capture at least part of the change, though the precision of boundary detection varies. In regions where there is a transition from bare land to road, our MFDCD and some others show better performance in accurately delineating the change area. Some methods, however, have varying degrees of missed detections. For instance, methods such as BIT [18] and ChangeFormer [19] seem to have less accurate delineation compared to our method and ELGC-Net [50]. In areas where buildings are under construction in one image and completed in another, different models have different performances. Some models, like ICIF [17] and TFI-GR [48], can detect the change but with somewhat blurred boundaries compared to the ground truth (white regions representing true positives). Other models may miss parts of the change. In areas where the change from bare land to road is affected by shadows (resulting in pseudo-changes), most models struggle. Except for our MFDCD, which shows relatively better performance in distinguishing true changes from pseudo-changes caused by shadows, other models either misclassify the shadow areas as changes (false positives in red) or miss the actual changes (false negatives in blue).
TABLE VIII: Ablation studies of DFC and TFF when trained on RB-SCD and public CD datasets. All values are reported as percentages (%). Results highlighted in red indicate the best performance, while those in blue denote the second-best performance.
Method DFC TFF RB-SCD LEVIR -CD SYSU -CD WHU -CD
(a) × × 50.45 81.65 68.98 86.65
(b) × 55.89 83.25 71.10 89.28
(c) × 55.23 82.97 70.89 88.50
(d) 57.52 84.69 72.63 90.43
TABLE IX: Effectiveness of the DFC module when trained on RB-SCD and public CD datasets. All values are reported in percentage (%). Results highlighted in red indicate the best performance, while those in blue denote the second-best performance.
Method RB-SCD LEVIR -CD SYSU -CD WHU -CD
(a) - 55.23 82.97 70.89 88.50
(b) $F_{i}^{1}+H_{i}^{2},i\in\left\{1,2,3,4\right\}$ 55.84 83.52 71.42 89.34
(c) $F_{i}^{t}+LL_{i}^{t},i\in\left\{1,2,3,4\right\}$ 56.10 83.67 71.68 89.45
(d) $F_{i}^{1}+H_{i}^{2},F_{i}^{2}+H_{i}^{1},i\in\left\{1,2\right\}$ 55.70 83.39 71.23 88.90
(e) $F_{i}^{t}+LL_{i}^{t},i\in\left\{3,4\right\}$ 55.86 83.51 71.30 89.24
(f) ASFF 56.42 84.02 72.06 89.98
(g) BTFF 56.87 83.90 71.85 89.77
(h) ASFF+BTFF 57.52 84.69 72.63 90.43
TABLE X: Effectiveness of graph filters with different numbers $k$ in TFF. All values are reported in percentage (%). Results highlighted in red indicate the best performance, while those in blue denote the second-best performance.
Method $k$ RB-SCD LEVIR-CD SYSU-CD WHU-CD
(a) - 55.89 83.25 71.10 89.28
(b) 3 56.74 83.84 71.98 89.87
(c) 4 57.13 84.13 72.31 90.10
(d) 5 57.52 84.69 72.63 90.43
(e) 6 57.40 84.26 72.25 90.38
TABLE XI: Effectiveness of MFDCD with Different Backbones when Trained on Four CD Datasets. All Values are Reported in Percentage (%). Results highlighted in red indicate the best performance, while those in blue denote the second-best performance.
Backbone Params (M) FLOPs (G) RB -SCD LEVIR -CD SYSU -CD WHU -CD
ResNet18 13.75 11.32 55.53 83.18 71.35 89.04
ResNet34 22.80 16.54 56.31 83.78 72.10 89.73
ResNet50 103.78 49.02 57.52 84.69 72.63 90.43
V-E Ablation Study
To demonstrate the effectiveness of our proposed modules, we conduct extensive experiments on RB-SCD and three public datasets. As shown in Tab. VIII, when the DFC module and TFF module are added to the baseline model, the mIoU on RB-SCD dataset increased by 5.44% and 4.78%, respectively. Additionally, the IoU on the LEVIR-CD, SYSU-CD, and WHU-CD datasets improved by 1.6%/2.12%/2.63% and 1.32%/1.91%/1.85%, respectively. As shown in Tab. IX, we conduct detailed ablation studies on the ASFF and BTFF blocks in the DFC module. Initially, we experiment with different coupling methods for ordinary features and frequency domain features. By comparing rows (b) and (c) with row (h), we can observe that ASFF and BTFF effectively dynamically couple visual features with frequency domain features, achieving the best results. Comparing rows (d) and (f), it can be seen that combining the low-level features at time $T_{1}$ with the corresponding high-frequency features at time $T_{2}$, due to semantic mismatches, results in worse performance compared to our designed ASFF. Comparing rows (e) and (g), it is clear that our designed BTFF captures the global structural changes across time, outperforming the coupling of high-level features and low-frequency features at the same time point. As shown in Tab. X, we conduct ablation experiments on the number of layers in the frequency filter of TFF. As the value of $k$ increased, the model’s performance improved. However, when the number of filter layers reached six, the performance stabilized or slightly declined. Subsequent experiments are based on the optimal setting, $k=5$.
To investigate the impact of backbone networks on the performance of the MFDD model in the CD tasks, as shown in Tab.  XI, in terms of parameters and computational complexity, ResNet18 has the lowest cost (13.75M parameters, 11.32G FLOPs), while ResNet50 has the highest (103.78M parameters, 49.02G FLOPs). In terms of detection accuracy, for all four datasets, the deeper the backbone network, the higher the accuracy. However, a deeper network also leads to a doubling of parameters and computational complexity. In this paper, ResNet50 is selected as the backbone network.
To more clearly illustrate the influence of our proposed modules, we visualized heatmaps for these modules. As depicted in columns (b)/(h) and (c)/(i) of Fig. 9, the low-level and high-level features from the backbone network typically exhibit certain characteristics. The low-level features may contain fine-grained but scattered information, while the high-level features, though more abstract, might lose some detailed spatial information. In columns (d)/(j), (e)/(k), and (f)/(l), after undergoing the ASFF block, BTTF block and TFF module respectively, the feature maps show progressive changes. The ASFF block helps to fuse features in a reasonable way, making the information more integrated. The BTTF block further refines the features, enhancing the relevant change information. The TFF module finally outputs feature maps that can better highlight the change regions, eliminating some irrelevant noise and enabling the acquisition of more distinct change-related information, which showcases the efficacy of the introduced modules in feature processing for CD tasks.
VI Conclusion
In this work, we introduce RB-SCD dataset, a SCD dataset specifically designed for complex traffic scenarios. The dataset features fine-grained pixel-level annotations across 11 semantic change categories and one background class, encompassing a wide range of road and bridge types. It also covers diverse geographic regions across multiple countries and urban environments, making it become a comprehensive and challenging benchmarks to date for road and bridge change analysis in RS imagery. Notably, RB-SCD supports both BCD and SCD tasks, offering flexibility for different application scenarios. By incorporating both common and rare semantic transformations, RB-SCD provides a valuable resource for the development and evaluation of CD models in traffic contexts.
To address the unique challenges presented by this dataset, we propose a novel multimodal CD framework, MFDCD, which leverages both spatial and frequency domain representations to enhance feature learning. By introducing frequency-domain analysis into the CD pipeline, MFDCD captures subtle semantic differences and structural changes that may be overlooked by conventional spatial-domain approaches. Our method also incorporates an effective fusion mechanism to integrate multimodal features, enabling more accurate detection of complex changes such as occluded road regions. Extensive experiments conducted on the RB-SCD dataset and public CD datasets demonstrate the superior performance of MFDCD across various metrics when compared to state-of-the-art CD methods. The results highlight the robustness, generalization ability, and semantic discrimination power of our approach in both straightforward and highly ambiguous change scenarios.
In future work, we plan to extend our exploration of road and bridge change detection by integrating additional modalities such as elevation data or temporal sequences, and further investigating the role of frequency-domain representations in improving the interpretability and adaptability of CD models. As the label imbalance analysis and challenges for RB-SCD dataset discussed in Sec. III-B, we also plan to expand the scale of the dataset and enrich it with more detailed spatiotemporal annotations to support broader research tasks such as urban infrastructure evolution and intelligent transportation systems. Furthermore, while MFDCD achieves strong performance, it is not yet lightweight enough for deployment in resource-constrained environments. Future research will focus on model compression and efficient architecture design to improve its applicability in real-time and edge scenarios.